OpenStudy (anonymous):

What is the slope of a line perpendicular to the line with equation y = 4x + 5? A. -1/4 B. 4 C. –4 D. –5

OpenStudy (anonymous):

D.

OpenStudy (anonymous):

ok are you sure

OpenStudy (anonymous):

It is not D

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

a perpendicular line will have a negative reciprocal slope. y = 4x + 5...the slope in this equation is 4. But we need the negative reciprocal of that slope. All that means is " flip " the slope and change the sign. For example : you are given slope 1/6. You want a perpendicular line slope, so you " flip " the slope and change the sign. So the slope you would need would be -6/1 (see how I flipped the slope and changed the sign) So what is the negative reciprocal of 4 ? Flip the slope and change the sign

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

4 is the same as 4/1. So " flip " the slope and change the sign. Not -4. Flip 4/1 and what do you get ? Then change the sign.

OpenStudy (anonymous):

1/4 so a

OpenStudy (anonymous):

or -1/4 oops sorry

OpenStudy (anonymous):

It is a, but do not forget to change the sign. It was positive...change it to negative. answer is -1/4. Understand ?

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

cool...that is always good to hear :)

OpenStudy (anonymous):

What is the slope of a line parallel to the line with equation 2x – 5y = 9? so same thing right let me try and maybe you can check i wanna make sure i pass

OpenStudy (anonymous):

you try and I will check. Keep in mind that a parallel line will have the same slope

OpenStudy (anonymous):

theres no choices you have to figure it yourself

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

put your equation in y = mx + b form, and the number in the m position is the slope

OpenStudy (anonymous):

2 is the slope in this equation

OpenStudy (anonymous):

no no 9

OpenStudy (anonymous):

2x - 5y = 9 -5y = -2x + 9 y = -2/-5x - 9/5 y = 2/5x - 9/5 slope is 2/5. Parallel line will have the same slope.

OpenStudy (anonymous):

do you understand or are you totally lost ?

OpenStudy (anonymous):

i do understand it better

OpenStudy (anonymous):

just remember, that in y = mx + b form, the slope is in the m position. And a parallel line will have the same slope. A perpendicular line will have negative reciprocal slope (thats when you flip the slope and change the sign)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

alright i will